In Math when you are dealing with Integers and Division, how do you divide with Like and Unlike signs?
Answer by cheesey(22)

Remember division is the opposite of multiplication. Since in division, your quotient multiplied by the divisor yields the dividend in question, look at the problem in terms of multiplication. If 20/2 = 10, (20 is the dividend and 2 is the divisor and 10 is the quotient), then 10 x 2 = 20. Do the same for signs, -20/2 = ? Since 2 is positive, how does one get a negative number (think multiplication; + x - = -), so the answer must be negative. Thus -20/2 = -10, since 2 x -10 = -20.
Thus remember, if the signs are different, your quotient is always negative. If the signs are the same, the quotient is always positive.
Hope this is helpful.